# Experimental Analysis of Biasing Voltage Generation in Wave-Controlled RIS

### v1

@@ I Introduction
The implementation of Reconfigurable Intelligent Surface (RIS) technology presents a novel way to improve wireless communication performance by providing adaptive degrees of freedom to design radio-frequency (RF) channels. Using an RIS, signals can be steered around obstacles, beamforming gains can be increased to boost signal-to-noise-ratio (SNR) and lower interference, and the overall quality-of-service (QoS) experienced by network users can be greatly enhanced [1]. RIS technology is especially beneficial in challenging RF environments like urban areas and indoor spaces, where it can be difficult to sustain strong and reliable signal transmissions [2].
@@
This paper investigates the performance of the technique presented in [13] and later studied in [14, 15, 16, 17] as a realistic and effective way to control RIS elements, as shown in Fig. 1. This technique, utilizing varactors for the reconfigurable reflective control, consists of delivering dc biasing voltages to the varactors inside the unit cells that are connected at specific positions, $x_{m}$, along a single biasing transmission line (BTL). The desired dc biasing voltages supplied to every RIS element to achieve a specific reflective radiation pattern, $w(x_{m})$, can be obtained from combinations of $N$ biasing standing waves (BSWs), $w(x_{m},t)$, that are injected into the BTL. The scope of this paper is limited to the study of the performance of the BTL when only a single frequency is injected into the BTL. Further, the extent to which a single reflected beam can be controlled through this single biasing frequency is explored. This technique drastically lowers signaling overhead and hardware complexity by doing away with the requirement for separate wired connections to every unit cell. Advanced RIS functions, including beam creation, polarization control, and channel equalization, are made possible by this straightforward yet efficient control method, which makes it a viable option for large-scale RIS deployments.
Figure 1: Wave-Controlled RIS System: The top layer consists of $M$ RIS elements, which are metallic patches connected to varactors uniformly spaced by $d_{x}$ along the $x$-direction in each row. The bottom layer contains NN biasing standing waves $w_{n}(x,t)$ that are excited on each BTL to generate the biasing voltages. These voltages are sampled at adjacent dc voltage outputs uniformly spaced by $d_{x}$ in the $x$-direction. Each row is controlled through only one connection on the right, where the NN frequencies are injected.
@@ II Biasing Standing Waves
The wave-controlled RIS approach involves two key components: a BTL board and an RIS board, as shown in Fig. 1. The BTL board, placed beneath the reflective surface and shielded from RF interference by a ground plane, supports $N$ BSWs. Rectifiers placed at $M$ positions on the BTL board generate dc voltages, $w(x_{m})$, from these waves which are used to bias the $M$ varactors positioned on the RIS board at $x_{m}$ for $m=0,1,\ldots,M-1$. This process modifies the reflected phase distribution, enabling control of the RIS’s scattered radiation pattern [16]. The necessary voltage applied to each varactor, $w(x_{m})$, for a given reflected phase is calculated using an analytical model that accounts for the RIS’s reflection behavior under various dc bias conditions for the varactors [18, 16]. Once the required voltages are determined, they are generated by a periodic time-domain signal comprising a combination of $N$ BSWs with specified amplitudes $W_{n}$, for $n=1,2,\ldots,N$, that are injected into the BTL through a single input port, along with a base voltage $W_{0}$. The inclusion of $W_{0}$ simplifies the reconstruction of the voltage distribution, as the varactor’s biasing voltage range typically does not include 0 V. Consequently, the expected voltage distribution pattern includes a dc component.
@@
As seen on the RIS board in Fig. 1, the RIS unit cells are separated by a distance $d_{x}$ along the $x$-axis. Likewise, the corresponding rectifiers that convert the standing wave amplitudes, $w_{n}(x_{m},t)$, into dc values for each RIS unit cell are also placed a distance $d_{x}$ away from each other along the $x$-axis on the BTL board. These, along with the voltage offset $W_{0}$, produce the necessary voltages $w(x_{m})$ [16]. The position along the BTL of the $m$th rectifier element is consequently defined as $x_{m}=md_{x}$.
@@
The BTL board is not necessarily as simple as that shown in Fig. 1, an example of which is provided in [15]. This BTL must connect with the RIS board at specific positions along the $x$-axis contained within the length $L$, where $L=(M-1)d_{x}$. The total length of the BTL along the $x$-axis is thus defined as $L_{\mathrm{tot}}=L+L_{\mathrm{left}}+L_{\mathrm{right}}$, where $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ are additional segments on either end of the BTL that help refine the voltage distribution pattern along the $M$ elements. The phase velocity of the propagated modes projected along the $x$-axis is $v_{\mathrm{b}}=c/n_{\mathrm{slow}}$, where $c$ represents the speed of light and $n_{\mathrm{slow}}\geq 1$ is a slowness factor determined by the material properties and geometry of the BTL. The injected periodic time-domain signal oscillates at the biasing frequency $f_{\mathrm{b}}=v_{\mathrm{b}}/\lambda_{\mathrm{b}}$, where $\lambda_{\mathrm{b}}=2\pi/k_{\mathrm{b}}$ is the guided wavelength, and the corresponding propagation wavenumber is $k_{\mathrm{b}}=2\pi f_{\mathrm{b}}/v_{\mathrm{b}}$.
@@
This design approach enables modeling of the voltage of the BSW before and after each rectifier. The voltage of the standing wave at any point along the BTL can be derived by applying transmission line theory [19]. In the general case when there are $N$ biasing modes (degrees of freedom to control the RIS response), the biasing voltage of the standing wave before rectification and assuming a short circuit termination at $x=-L_{\mathrm{left}}$ is [13, 16]
@@
Here, $n$ is the biasing mode index and $W_{n}$ is the real-valued amplitude of the standing wave of the $n$th biasing mode. The term $W_{0}$ is the dc component, which sets the minimum biasing voltage applied to the BTL and shifts the varactor’s operating range. Furthermore, $\phi_{n}$ is the temporal phase offset of the $n$th biasing mode. This definition generalizes Eq. (7) of Ref. [16], as here the wavelength $\lambda_{\mathrm{b}}$ and, consequently, the fundamental frequency $f_{\mathrm{b}}$, do not necessarily depend on the total length of the BTL $L_{\mathrm{tot}}$.
@@
$\begin{split}&w(x_{m})=W_{0}+\\ &\max_{t}\left[\sum_{n=1}^{N}W_{n}\sin\left(\frac{n\omega_{\mathrm{b}}n_{\mathrm{slow}}}{c}\left(x_{m}+L_{\mathrm{left}}\right)\right)\cos(n\omega_{\mathrm{b}}t+\phi_{n})\right].\end{split}$ (2)
@@
$w(x_{m})=W_{0}+\left|W_{\mathrm{b}}\sin\left(\frac{\omega_{\mathrm{b}}n_{\mathrm{slow}}}{c}\left(x_{m}+L_{\mathrm{left}}\right)\right)\right|.$ (4)
@@ III System Design
The wave-controlled RIS system is realized here using two separate boards: the RIS board and the BTL board. The RIS board reflects the incident radiation based on a spatial distribution of the phase reflection coefficient across the RIS elements, which is governed by the biasing voltage distribution $w(x_{m})$ provided by the BSW, supported by the BTL board. Other implementations could be realized in a single multilayer board. This biasing pattern is controlled by the signal frequencies and amplitudes injected into the BTL (see Fig. 1). The RIS unit cell is designed to achieve the desired phase response by applying a variable voltage to the varactor. The design includes the unit cell’s geometry, material, and chosen varactor. The BTL board comprises a transmission line with a defined geometry and integrated rectifiers to provide the dc biasing voltage for the unit cells of the RIS. Material specifications play a crucial role in the overall performance of the BTL and are introduced next.
@@ III-A RIS Unit Cell Design
The RIS unit cell consists of an array of two rectangular patches arranged in a mirrored configuration as in Fig. 2, mounted on a grounded dielectric substrate (RT5880LZ) with a relative permittivity of $\epsilon_{r}=2$, a loss tangent of $\tan\delta=0.0021$, and a thickness of 1.27 mm. Each patch is connected to the cathode of an individual SMV1231-040LF varactor, while the varactor’s anode is grounded via a through-hole connection. The unit cell design is optimized to maximize the phase dynamic range within the industrial, scientific, and medical (ISM) 2.45 GHz band for reverse bias voltages ranging from 4-4 V to approximately 10-10 V.
To interface the patch with the BTL board and allow the dc biasing of the two varactors, a coplanar waveguide with ground is implemented on the ground plane of the RIS and connected to the rectified voltage $w(x_{m})$ in the BTL. This line is oriented to avoid disrupting the RIS ground-plane current and is connected to the center of the patch through a metallic via. This configuration allows efficient biasing of the varactors while minimizing the effect on the RIS RF current flow. To further suppress the effect on the RF current at the connection node located at the center of the coplanar waveguide, a 100 pF capacitor is incorporated to ensure proper grounding at RF. This capacitor is also part of the rectifier design that is discussed in Sec. III B. The RIS board comprises 27 identical unit cells with two copper patches, with each unit cell featuring the following dimensions (in mm): $d_{x}=20$, $A_{1}=18.5$, $A_{2}=1$, $A_{3}=0.3$, $B=145.4$, $B_{1}=36.7$, $B_{2}=1.2$, $B_{3}=24.8$, $G_{1}=0.4$, $G_{2}=0.5$, $G_{3}=1.8$, $r=0.2$, $P_{1}=4$, $P_{2}=0.4$, as shown in Fig. 2.
@@
To calculate the reflected phase provided by each RIS element, the analytical circuit model shown in Fig. 3 is used [18, 16]. This model requires only one full-wave simulation of the unit cell over the frequency of operation (assuming the varactor is disconnected) to extract the circuit elements that represent just the patch over the grounded slab, as detailed in Appendix A. For the design presented in Fig. 2, after fine-tuning, the values obtained are $R_{\mathrm{d}}=0.17$ $\Omega$, $C_{\mathrm{d}}=0.74$ pF, $L_{\mathrm{d}}=1.64$ nH, and $L_{\mathrm{s}}=1.60$ nH.
Once these circuit elements are determined, the varactor is incorporated to complete the RIS unit cell model by adding its RF impedance $Z_{\mathrm{v}}$ to the equivalent circuit as shown in Fig. 3. The varactor’s impedance is modeled by an equivalent series RLC circuit
$Z_{\mathrm{v}}=R_{\mathrm{v}}(V)+j\omega L_{\mathrm{v}}+\frac{1}{j\omega C_{\mathrm{v}}(V)},$ (5)
where the variable capacitance $C_{\mathrm{v}}\mathrm{(V)}$ and variable resistance $R_{\mathrm{v}}\mathrm{(V)}$ are functions of the reverse biasing voltage $V$, and the constant inductance $L_{\mathrm{v}}=2.34$ nH models both the package inductance and the parasitic inductance introduced when the varactor is soldered across the unit cell gap. The values of $C_{\mathrm{v}}\mathrm{(V)}$ and $R_{\mathrm{v}}\mathrm{(V)}$ are obtained through a parametric sweep simulation in the commercial Keysight Advanced Design System (ADS) software, based on the SPICE model of the varactor provided by the manufacturer. These values are limited to the reverse voltage bias range of $[4$ V, $15$ V]] and are detailed in Table I.
@@
$Z_{\text{RIS}}=\left(R_{\mathrm{d}}+j\omega L_{\mathrm{d}}+Z_{\mathrm{v}}\parallel\frac{1}{j\omega C_{\mathrm{d}}}\right)\parallel j\omega L_{\mathrm{s}}.$ (6)
Finally, $Z_{\text{RIS}}$ is used to determine the reflection coefficients, $\Gamma_{m}$, of the RIS elements:
$\Gamma_{m}=\frac{Z_{\mathrm{RIS}}-\eta_{0}}{Z_{\mathrm{RIS}}+\eta_{0}},$ (7)
where $\eta_{0}=377$ $\Omega$ is the free-space impedance. The magnitude and phase of $\Gamma_{m}$ are denoted by $R_{m}(\mathrm{V})$ and $\alpha_{m}(\mathrm{V})$, respectively. The values of $R_{m}$ and $\alpha_{m}$ as functions of the biasing voltage for the unit cell in Fig. 2 are shown in Fig. 4(4(a)) for three different frequencies, including the target RF frequency $f_{\mathrm{c}}=2.45$ GHz. The voltage $\mathrm{V}$ is positive as the varactor is connected in reverse, with its anode grounded and its cathode connected to the positive biasing voltage. At $f_{\mathrm{c}}=2.45$ GHz, the phase $\alpha_{m}$ varies between $160^{\circ}$ and $140^{\circ}$, with the most significant variations occurring between $4$ and $10$ V. The lowest reflection magnitude $R_{m}$, approximately 2.4-2.4 dB, is observed around $7$ V. Fig. 4(4(b)) illustrates the behavior of $\alpha_{m}$ and $R_{m}$ as functions of the RF frequency and compares the analytical model (solid lines) with full-wave simulations performed using the commercial software CST Studio Suite (dashed lines) for five different biasing voltages. The agreement is generally good across different biasing voltages and frequencies, with particularly strong accuracy at the target frequency $f_{\mathrm{c}}=2.45$ GHz (dotted line in blue). Although in practical RIS implementations the control is optimized based on channel estimates, having an analytical model is useful for design and research purposes or to assist the optimization algorithms.
Figure 4: Phase and magnitude response of the reflection coefficient for the RIS unit cell as a function of (a) the varactor’s biasing voltage at three different frequencies, based on the analytical model described by (7), and (b) the frequency at five different biasing voltages. Solid lines represent results from (7) and dashed lines correspond to full-wave simulations.

### v2

@@ I Introduction
The implementation of reconfigurable intelligent surface (RIS) technology presents a novel way to improve wireless communication performance by providing adaptive degrees of freedom to design radio-frequency (RF) channels. Using an RIS, signals can be steered around obstacles, beamforming gains can be increased to boost signal-to-noise ratio (SNR) and lower interference, and the overall quality of service (QoS) experienced by network users can be greatly enhanced [1]. RIS technology is especially beneficial in challenging RF environments like urban areas and indoor spaces, where it can be difficult to sustain strong and reliable signal transmissions [2].
@@
This paper investigates the performance of the technique presented in [13] and later studied in [14, 15, 16, 17, 18] as a realistic and effective way to control RIS elements. A schematic is shown in Fig. 1. This technique, utilizing varactors for the reconfigurable reflective control, consists of delivering dc biasing voltages to the varactors inside the unit cells that are connected at specific positions, $x_{m}$, along a single biasing transmission line (BTL). The desired dc biasing voltages supplied to every RIS element to achieve a specific reflective radiation pattern, $w(m)$, can be obtained from combinations of $N$ biasing standing waves (BSWs), $w(x_{m},t)$, that are injected into the BTL. The scope of this paper is limited to the study of the performance of the BTL when only a single frequency is injected into the BTL. Further, the extent to which a single reflected beam can be controlled through this single biasing frequency is explored. This technique drastically lowers signaling overhead and hardware complexity by doing away with the requirement for separate wired connections to every unit cell. Advanced RIS functions, including beam creation, polarization control, and channel equalization, are made possible by this straightforward yet efficient control method, which makes it a viable option for large-scale RIS deployments.
Figure 1: Wave-Controlled RIS System: The top layer consists of $M$ RIS elements, which are metallic patches connected to varactors uniformly spaced by $d_{x}$ along the $x$ direction in each row. The bottom layer contains NN biasing standing waves $w_{n}(x,t)$ that are excited on each BTL. Rectifiers uniformly spaced by $d_{x}$ in the $x$ direction are used to generate the dc biasing voltages for the varactors. Each BTL is controlled through only one connection on the right, where the NN frequencies are injected. Here, a BTL is straight just for conceptualization, though in practice we use a meandered line to be able to use very low-frequency BSWs.
@@ II Biasing Standing Waves
The wave-controlled RIS approach involves two key components: a BTL board and an RIS board, as shown in Fig. 1. The BTL board, placed beneath the reflective surface and shielded from RF interference by a ground plane, supports $N$ BSWs. Rectifiers placed at $M$ positions on the BTL board generate dc voltages, $w(m)$, from these waves which are used to bias the $M$ varactors positioned on the RIS board at $x_{m}$ for $m=0,1,\ldots,M-1$. This process modifies the reflected phase distribution, enabling control of the RIS’s scattered radiation pattern [16]. The necessary voltage applied to each varactor, $w(m)$, for a given reflected phase is calculated using an analytical model that accounts for the RIS’s reflection behavior under various dc bias conditions for the varactors [19, 16]. Once the required voltages are determined, they are generated by a periodic time-domain signal comprising a combination of $N$ BSWs with specified amplitudes $W_{n}$, for $n=1,2,\ldots,N$, that are injected into the BTL through a single input port, along with a base voltage $W_{0}$. The inclusion of $W_{0}$ simplifies the reconstruction of the voltage distribution, as the varactor’s biasing voltage range typically does not include 0 V. Consequently, the expected voltage distribution pattern includes a dc component.
@@
As seen on the RIS board in Fig. 1, the RIS unit cells are separated by a distance $d_{x}$ along the $x$ direction. Likewise, the corresponding rectifiers that convert the standing wave amplitudes, $w_{n}(x_{m},t)$, into dc values for each RIS unit cell are also placed a distance $d_{x}$ away from each other along the $x$ direction on the BTL board. These, along with the voltage offset $W_{0}$, produce the necessary voltages $w(m)$ [16]. The position along the BTL of the $m$th rectifier element is consequently defined as $x_{m}=md_{x}$.
@@
The BTL board is not necessarily as simple as that shown in Fig. 1, an example of which is provided in [15]. This BTL must connect with the RIS board at specific positions along the $x$ direction within the length $L$, where $L=(M-1)d_{x}$. The total length of the BTL along the $x$ direction is thus defined as $L_{\mathrm{tot}}=L+L_{\mathrm{left}}+L_{\mathrm{right}}$, where $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ are additional segments on either end of the BTL that help refine the voltage distribution pattern along the $M$ elements. The phase velocity of the propagated modes projected along the $x$ direction is $v_{\mathrm{b}}=c/n_{\mathrm{slow}}$, where $c$ represents the speed of light and $n_{\mathrm{slow}}\geq 1$ is a slowness factor determined by the material properties and geometry of the BTL. The injected periodic time-domain signal oscillates at the biasing frequency $f_{\mathrm{b}}=v_{\mathrm{b}}/\lambda_{\mathrm{b}}$, where $\lambda_{\mathrm{b}}=2\pi/k_{\mathrm{b}}$ is the guided wavelength, and the corresponding propagation wavenumber is $k_{\mathrm{b}}=2\pi f_{\mathrm{b}}/v_{\mathrm{b}}$.
@@
This design approach enables modeling of the voltage of the BSW before and after each rectifier. The voltage of the standing wave at any point along the BTL can be derived by applying transmission line theory [20]. In the general case when there are $N$ biasing modes (degrees of freedom to control the RIS response), the biasing voltage of the standing wave before rectification and assuming a short circuit termination at $x=-L_{\mathrm{left}}$ is [13, 16, 18]
@@
Here, $n$ is the biasing mode index and $W_{n}$ is the real-valued amplitude of the standing wave of the $n$th biasing mode. The term $W_{0}$ is the dc component, which sets the minimum biasing voltage applied to the BTL and shifts the varactor’s operating range. Furthermore, $\phi_{n}$ is the phase offset of the $n$th biasing mode. This definition generalizes Eq. (7) of Ref. [16], as here the wavelength $\lambda_{\mathrm{b}}$ and, consequently, the fundamental frequency $f_{\mathrm{b}}$, do not necessarily depend on the total length of the BTL $L_{\mathrm{tot}}$; in other words, we do not need standing waves that are resonant with the BTL length.
@@
$\begin{split}&w(m)=W_{0}+\\ &\max_{t}\left[\sum_{n=1}^{N}W_{n}\sin\left(\frac{n\omega_{\mathrm{b}}n_{\mathrm{slow}}}{c}\left(x_{m}+L_{\mathrm{left}}\right)\right)\cos(n\omega_{\mathrm{b}}t+\phi_{n})\right].\end{split}$ (2)
@@
$w(m)=W_{0}+\left|W_{\mathrm{b}}\sin\left(\frac{\omega_{\mathrm{b}}n_{\mathrm{slow}}}{c}\left(x_{m}+L_{\mathrm{left}}\right)\right)\right|.$ (4)
@@ III System Design
The wave-controlled RIS system is realized here using two separate boards: the RIS board and the BTL board. The RIS board reflects the incident radiation based on a spatial distribution of the phase reflection coefficient across the RIS elements, which is governed by the biasing voltage distribution $w(m)$ provided by the BSW, supported by the BTL board. Other implementations could be realized in a single multilayer board. This biasing pattern is controlled by the signal frequencies and amplitudes injected into the BTL (see Fig. 1). The RIS unit cell is designed to achieve the desired phase response by applying a variable voltage to the varactor. The design includes the unit cell’s geometry, material, and chosen varactor. The BTL board comprises a transmission line with a defined geometry and integrated rectifiers to provide the dc biasing voltage for the unit cells of the RIS. Material specifications play a crucial role in the overall performance of the BTL and are introduced next.
@@ III-A RIS Unit Cell Design
The RIS unit cell consists of an array of two rectangular patches arranged in a mirrored configuration as in Fig. 2, mounted on a grounded dielectric substrate (RT5880LZ) with a relative permittivity of $\epsilon_{r}=2$, a loss tangent of $\tan\delta=0.0021$, and a thickness of 1.27 mm. Each patch is connected to the cathode of an individual SMV1231-040LF varactor, while the varactor’s anode is grounded via a through-hole connection. We assume an incident $y$-polarized electric field, and the unit cell design is optimized to maximize the phase dynamic range within the industrial, scientific, and medical (ISM) 2.45 GHz band for reverse bias voltages ranging from $4$ V to approximately $10$ V.
The rectified reverse voltage $w(m)$ is provided to a varactor on the top RIS layer in the $m$th unit cell through a vertical biasing via from the center of the RIS patch to a coplanar waveguide on the ground plane of the RIS board (Fig. 2). This line is oriented to avoid disrupting the RIS ground-plane current. This configuration allows efficient biasing of the varactors while minimizing the effect on the RIS RF current flow. To further suppress the effect on the RF current at the connection node located at the center of the coplanar waveguide, the capacitor $C_{r1}=100$ pF is incorporated to ensure proper grounding at RF (Fig. 2). This capacitor is also part of the rectifier in the BTL board, as discussed in Sec. III B. The RIS board comprises 27 identical unit cells with two copper patches, with each unit cell featuring the following dimensions (in mm): $d_{x}=20$, $A_{1}=18.5$, $A_{2}=1$, $A_{3}=0.3$, $B=145.4$, $B_{1}=36.7$, $B_{2}=1.2$, $B_{3}=24.8$, $G_{1}=0.4$, $G_{2}=0.5$, $G_{3}=1.8$, $r=0.2$, $P_{1}=4$, $P_{2}=0.4$, as shown in Fig. 2.
@@
To calculate the reflected phase provided by each RIS element, the analytical circuit model shown in Fig. 3 is summarized in [19, 16, 18]. This model requires only one full-wave simulation of the unit cell over the frequency of operation (assuming the varactor is disconnected) to extract the circuit elements that represent just the patch over the grounded slab, as detailed in Appendix A. For the design presented in Fig. 2, the values obtained are $L_{\mathrm{s}}=1.60$ nH, $R_{\mathrm{d}}=0.17$ $\Omega$, $L_{\mathrm{d}}=1.92$ nH, and $C_{\mathrm{d}}=0.98$ pF represents the two fringe capacitances at the two sides of each patch toward the ground and toward adjacent patch elements, as shown in Fig. 3 of [21] and Fig.13 of [22].
Once these circuit elements are determined, the varactor is incorporated to complete the RIS unit cell model by adding its RF impedance $Z_{\mathrm{v}}$ to the equivalent circuit as shown in Fig. 3. In this configuration, the varactor is physically connected only on one side of the patch toward the ground, and therefore its impedance $Z_{\mathrm{v}}$ is in parallel to only one of the two fringe capacitors. Accordingly, the lumped capacitor $C_{\mathrm{d}}$ is divided into two series capacitors and, after tuning, the fitting with full-wave results provides the values $C_{\mathrm{d1}}=1.25$ pF and $C_{\mathrm{d2}}=2.72$ pF. We have verified that the model in Fig. 3 with split capacitors $C_{\mathrm{d1}}$ and $C_{\mathrm{d2}}$ produces more accurate results than considering a single capacitor $C_{\mathrm{d}}$ in parallel with the varactor impedance, over the whole range of frequencies and varactor voltages considered. The circuit in Fig. 3 replaces Fig. 4 in [18], where the splitting of the capacitance $C_{\text{d}}$ into two separate capacitors, $C_{\text{d1}}$ and $C_{\text{d2}}$, was omitted from the figure but accounted for in the numerical results.
The varactor’s impedance is modeled by an equivalent series RLC circuit
$Z_{\mathrm{v}}(V)=R_{\mathrm{v}}(V)+j\omega L_{\mathrm{v}}+\frac{1}{j\omega C_{\mathrm{v}}(V)},$ (5)
where the variable capacitance $C_{\mathrm{v}}\mathrm{(V)}$ and variable resistance $R_{\mathrm{v}}\mathrm{(V)}$ are functions of the reverse biasing voltage $V$, and the constant inductance $L_{\mathrm{v}}=2.39$ nH models both the package inductance and the parasitic inductance introduced when the varactor is soldered across the unit cell gap. The values of $C_{\mathrm{v}}\mathrm{(V)}$ and $R_{\mathrm{v}}\mathrm{(V)}$ are obtained through a parametric sweep simulation in the commercial Keysight Advanced Design System (ADS) software, based on the SPICE model of the varactor provided by the manufacturer. These values are limited to the reverse voltage bias range of $[4$ V, $15$ V]] and are detailed in Table I.
@@
$Z_{\text{RIS}}=\left(R_{\mathrm{d}}+j\omega L_{\mathrm{d}}+\frac{1}{j\omega C_{\mathrm{d2}}}+Z_{\mathrm{v}}\parallel\frac{1}{j\omega C_{\mathrm{d1}}}\right)\parallel j\omega L_{\mathrm{s}}.$ (6)
Finally, $Z_{\text{RIS}}$ is used to determine the reflection coefficients, $\Gamma$, of the RIS elements:
$\Gamma=\frac{Z_{\mathrm{RIS}}-\eta_{0}}{Z_{\mathrm{RIS}}+\eta_{0}},$ (7)
where $\eta_{0}=377$ $\Omega$ is the free-space impedance. The magnitude and phase of $\Gamma$ are denoted by $R(\mathrm{V})$ and $\alpha(\mathrm{V})$, respectively. The values of $R$ and $\alpha$ as functions of the biasing voltage for the unit cell in Fig. 2 are shown in Fig. 4(4(a)) for three different frequencies, including the target RF frequency $f_{\mathrm{c}}=2.45$ GHz. The voltage $\mathrm{V}$ is positive as the varactor is connected in reverse, with its anode grounded and its cathode connected to the positive biasing voltage. At $f_{\mathrm{c}}=2.45$ GHz, the phase $\alpha$ varies between $160^{\circ}$ and $140^{\circ}$, with the most significant variations occurring between $4$ and $10$ V. The lowest reflection magnitude $R$, approximately 2.4-2.4 dB, is observed around $7$ V. Fig. 4(4(b)) illustrates the behavior of $\alpha$ and $R$ as functions of the RF frequency and compares the analytical model (solid lines) with full-wave simulations performed using the commercial software CST Studio Suite (dashed lines) for five different biasing voltages. The agreement is generally good across different biasing voltages and frequencies, with particularly strong accuracy at the target frequency $f_{\mathrm{c}}=2.45$ GHz (dotted line in blue). Although in practical RIS implementations the control is optimized based on channel estimates, having an analytical model is useful for design and research purposes or to assist the optimization algorithms.
The model is obtained using periodic boundary conditions for normal incidence. As conventionally done in reflectarrays or RIS metasurfaces, we use the well accepted “local periodicty” approximation: if the $m$th varactor has a voltage bias $w(m)$ at $x_{m}=md_{x}$ for m=0,1,,M1m=0,1,...,M-1, that is slowly varying with mm, then the $m$th reflection coefficient $\Gamma_{m}=R_{m}e^{j\alpha_{m}}$ is well approximated by (7) .
Figure 4: Phase and magnitude response of the reflection coefficient for RIS unit cells as a function of (a) the varactor’s biasing voltage at three different frequencies, based on the analytical model described by (7), and (b) the frequency at five different biasing voltages. Solid lines represent results from (7) that are in good agreement with full-wave simulations (dashed lines).